### v1

@@ Abstract
Do social networks and peer influence shape major life decisions in highly polarized settings? We explore this question by examining how peers influenced the allegiances of West Point cadets during the American Civil War. Leveraging quasi-random variations in the proportion of cadets from Free States, we analyze how these differences affected decisions about which army to join. We find that a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army, while almost all cadets from Free States joined the Union Army (if they decided to join the war). We further examine how cadets’ decisions affected their military rank and career outcomes. Our findings highlight that peers still influence choices even when they are life-altering and occur during periods of extreme polarization.
@@ 1 Introduction
In a highly polarized world, ideological divides not only reflect personal opinions but also actively shape critical decisions, from health choices like vaccination uptake to alignment in global conflicts. Because of these divides, individuals often grapple with challenging decisions and conflicting incentives. For example, during the COVID-19 pandemic, focus groups and surveys suggest that social expectations—–sometimes at odds with personal preferences—–played a significant role in vaccination decisions (Lin et al., 2022). Similarly, in politically polarized settings, it has been suggested that decisions about voting or expressing political views can be deeply influenced by social networks, as individuals may suppress their private preferences to conform to publicly dominant views (Kuran, 1998). Theory has long suggested that social networks play a critical role in shaping these decisions, not only through peer pressure for conformity or allegiance, but also by providing insight and information and sharing relevant experiences, and providing comradery (e.g., Bernheim, 1994, Jackson, 2008).
Although there is extensive evidence of peer effects in a variety of settings (Benhabib, Bisin and Jackson, 2010; Brechwald and Prinstein, 2011; McGloin and Thomas, 2019; Giletta et al., 2021), it is unclear whether peers still sway decisions when other economic and social forces become enormous and a single decision can be life-altering. Providing causal evidence on how social networks influence behavior in highly polarized and life-changing environments remains challenging because individuals tend to associate with like-minded peers, making it difficult to disentangle the influence of social networks from preexisting shared beliefs and communal forces.
In this paper, we address this challenge by examining the historical context of the American Civil War, one of the most polarized periods in U.S. history. Specifically, we investigate how West Point cadets–—a group that later served as leaders in both the Union and Confederate armies–—were influenced by their peers when deciding their allegiances. Although West Point admissions broadly reflected the geographic distribution of congressional districts over time, annual recruitment patterns varied significantly due to the decentralized nomination process and qualification exams (e.g., Park, 1840). We show that this year-to-year variation in peer composition was quasi-random. Exploiting this feature, we provide evidence of the significant role that peers played in shaping decisions during a moment of extreme polarization, and where other background communal and geographic forces might have seemed otherwise overwhelming.
The American Civil War was a pivotal event in U.S. history, notable not only for its significant political and social impacts but also for the personal conflicts it caused among individuals torn between nationalism and sectionalism. West Point cadets, who later became leaders of both the Union and Confederate armies, were among those who faced split allegiances. Although these cadets played vital roles in the war, there has been limited empirical investigation into the factors influencing their decision on which side to fight for. [^1] There exists a rich historical literature on West Pointers in the Civil War (e.g., Patterson, 2002). A recent political science study, White (2024) documents the importance of economic interest in influencing cadets’ choices. See additional reference below.
West Point, as the premier military training institution of the time, offered a unique environment where cadets from both Northern (Free) and Southern (Slave) States lived, trained, and formed close-knit cohorts prior to the war. This setting makes peer influence a particularly relevant dynamic for study. There is research suggesting that peer influence and “group cohesion” played important roles in motivating people to enlist (and fight), but much of the evidence is anecdotal and drawn from letters and diaries (e.g., see McPherson (1997, 2003); Siebold (2007)). How interactions between cadets from these different regions influenced allegiance during the Civil War has yet to be explored with a careful statistical analysis. Did exposure to northern peers sway southern cadets toward the Union cause, or did the sociopolitical context of their home states dominate and insulate them from peer influence?
@@
We study these questions by digitizing detailed biographical data on over 1,600 West Point cadets from 1820 to 1860 and leveraging quasi-random variations in peer composition. We classify states as slave or free based on whether slavery was legal in 1860 (which corresponds to whether the slave population in the 1860 census exceeded 1%), and we revisit the results with alternative thresholds for robustness. The share of peers from Free States fluctuated annually without clear trends, providing variation in peer exposure (see Section 3.1 for institutional reasons that drive the variation). This variation is not systematically correlated with individual characteristics or home-state differences—whether considering the full sample or restricting to cadets who later participated in the war—consistent with the interpretation of peer composition as quasi-random. This quasi-randomness allows us to overcome common endogeneity concerns in identifying peer effects (e.g., Manski, 1993; Angrist, 2014). In the main text, we focus on the wartime choices of cadets who served in the conflict. The Supplemental Appendix presents additional robustness checks using the full cadet sample.
@@
Our main finding is that a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. The effect is sizable: a one-standard deviation increase in the proportion of peers from Free States raised the probability of joining the Union by four percentage points, or 11% of the mean. In contrast, nearly all cadets from Free States joined the Union regardless of cohort composition. The difference suggests a conflict between nationalism and sectionalism faced by cadets from Slave States, who were more susceptible to peer influence. Furthermore, within the Slave States, we find that peer influence mattered little for cadets from states in which the slave population was high (e.g., above 1/3), but mattered strongly for cadets from Slave States with slave populations below 1/3, suggesting an interaction between peer influence and economic considerations rooted in the slave economy. Indeed, the peer influence was large and almost entirely concentrated among cadets from Border and Mid-level Slave States where support for secession was more divided.
We present several additional analyses to examine heterogeneity and test the robustness of our results. Notably, peer effects are stronger when Slave States are defined using thresholds of 5% or 10% slave population, suggesting that states with minimal slave populations resemble Free States in their influence. Moreover, peer influence is more pronounced within the same cohort than across cohorts, indicating the importance of direct, contemporaneous interactions. Similarly, cadets who completed their training exerted greater influence than those who did not.
@@
Because many West Point cadets in our study later emerged as significant military figures, we further examine how their wartime choices influenced their post-war trajectories. Employing both OLS and an instrumental variable strategy (where home state and peer exposure predict allegiance), we find that for those from Slave States, those who sided with the Union experienced lower military ranks in 1865, a diminished chance of becoming generals, and a decreased likelihood of wartime mortality.
The variation in cohort composition used in our main analysis does not pinpoint the mechanism through which peer influence operates. To probe this further, we examine variation in cadets’ likelihood of having co-fought in the Mexican-American War (1846–1848). We find that peer effects were stronger when slave-state cadets had more free-state peers with whom they had shared combat experience, and weaker when their co-combatants were also from Slave States. Due to the limited sample—restricted to slave-state cadets who graduated before 1846—the interaction estimates are sizable but imprecise. While not conclusive, this pattern suggests that shared formative experiences, such as fighting side-by-side, played a role in peer influence.
Our study contributes to the broad social science literature on peer effects (see Benhabib, Bisin and Jackson (2010); Brechwald and Prinstein (2011); McGloin and Thomas (2019); Giletta et al. (2021) for some overviews), by demonstrating how peer influence can shape major life decisions, particularly during moments of extreme polarization. The experiences of West Point cadets during the American Civil War mirror other historical contexts where military academies produced leaders for both sides of major conflicts in divided nations. For instance, the Whampoa Military Academy in the 1920s produced commanders who fought on opposing sides during the Chinese Civil War (e.g., Bai, Jia and Wang, 2023), and the Heroic Military Academy in Mexico trained leaders who both supported and opposed the regime during the Mexican Revolution.
@@
As mentioned above, such decisions are not unique to military cadets; similar choices are faced by ordinary citizens in many divided societies. However, analyzing peer influence in these broader contexts is often more challenging. The distinctive setting of the American Civil War, where allegiances were closely tied to geography, provides a unique opportunity to examine peer influence with a level of identification that is difficult to achieve in other cases.
@@
Finally, our study contributes to the expanding literature on leaders and leadership (e.g., Jones and Olken, 2005; Acemoglu and Jackson, 2015; Dippel and Heblich, 2021; Cagé et al., 2023; Bai, Jia and Yang, 2023). Understanding how leaders develop their ideas and loyalties is central to grasping important political and economic outcomes. Our findings highlight the significance of peer influence in shaping future leaders’ decisions during crises, demonstrating how peer influence can drive individual choices that may shape the course of history.
@@ 2.2 Geography and the Classification of States by High/Low Slave
In our data, 964 and 674 cadets came from Free States and Slave States, respectively. Among those from Free States, 540 joined the armies and 91.9% of them joined the Union. Among those from Slave States, 389 joined the armies and 36.2% of them joined the Union. Supplemental Appendix Table S2 presents the summary statistics of the key variables using the baseline threshold.
@@ 3.2 Statistical Analysis
$Union_{i,s,t}=\beta_{0}+\beta_{1}Peer_{i,t}+\gamma X_{i,s}+\alpha_{s}+t+\epsilon_{i,s,t},$ (1)
where $Union_{i,s,t}$ denotes whether cadet $i$ from state $s$ and cohort $t$ joined the Union. $Peer_{i,t}$ represents the proportion of peers from Free States. $\alpha_{s}$ represents state fixed effects, whereas $X_{i,s}$ includes a range of personal and state characteristics such as the cadet’s age in 1860, their academic rank at West Point and the proportion of the slave population in their home state. Naturally, state characteristics are absorbed once home state fixed effects are accounted for. We incorporate a linear time trend $t$ to account for potential temporal trends in Union enlistment, as there is a slight upward trend when considering our entire time horizon. To account for potential serial correlations, we employ bootstrapped standard errors with 400 resampling iterations. We also estimate the model using standard errors clustered at the cohort level, and the statistical significances are unchanged.
@@ 3.3 The Impact of Peers on Cadets’ Allegiance
Table 1 displays our main baseline results, indicating that peer influence matters for individuals from Slave States. Column (1) illustrates the raw correlation between the proportion of Free-State peers and the likelihood of joining the Union. Column (2) incorporates personal and state characteristics, while Column (3) further includes state fixed effects. To facilitate the interpretation of coefficients, we present the coefficient for a one standard deviation change in peer exposure. As shown, these estimates, approximately 0.04, are consistent across different models, confirming the quasi-randomness of peer composition. The effect is sizable, representing about 11% of the mean of the dependent variable (0.36).111111 This is a conservative estimate. When using alternative thresholds (5% and 10%), we find that peer effect raises the probability of Slave-State cadets joining the union by 17-19% (see Supplemental Appendix Table S4).
@@
Relevance of Co-fighting in the Mexican-American War
3.4 Additional Analyses
We conduct six supplementary analyses to explore (1) whether peer influence within a shared cohort is of greatest significance (compared to adjacent cohorts), (2) how the effect of peers on cadets varied across different various time-frames, (3) potential differences between peers who graduated and those who dropped out, (4) how incorporating non-participants of the war affects our conclusions, (5) how accounting for cash-crop employment shapes peer influence, and (6) estimates based on defining peers by appointment state rather than home state. We describe the results from these analyses below.
Influence by Cohort
We analyze peers by grouping them into cohorts and display the coefficients for peers within the same cohort, as well as for cohorts from t-1 to t-2 and t+1 to t+2, in Supplemental Appendix Figure S3. The coefficients in the figure are obtained from a single regression that includes all controls from our baseline analysis. The result reveals that peer influence is most important within the same cohort. In contrast, although there is some degree of influence from earlier cohorts, future cohorts do not show comparable levels of influence.
Peer Influence pre-post 1850
Although the issue of slavery was highly contentious throught the period we study, it intensified in the 1850sand nearly led to the collapse of the United States. In an effort to avert civil war, Congress passed the Compromise of 1850, which included the Fugitive Slave Law that obligated law enforcement nationwide to assist in capturing alleged runaway slaves, and many in the north resisted its enforcement (McPherson, 2003). Against this backdrop, we use 1850 as a dividing point to assess whether peer influence experienced notable changes before and after that year. As illustrated in Supplemental Appendix Table S7, peer influence (though estimated imprecisely due to limited sample sizes) persisted over time, with a similar effects on either side of this divide.
Graduate vs. Dropout Peer Influence
Including Those Who Were Not in the Armies
Considering Cash-Crop Employment
Considering Appointment State
@@ 3.5 Career Outcomes
We know the post-war outcomes of many who joined an army, allowing us to investigate how their decisions during this critical historical juncture influenced their life results, such as military rank and survival probability.
As indicated in Columns (1)-(2) of Table 3, joining the Union correlates with a lower military rank in 1865 and a reduced probability of attaining a rank of General-level positions. This aligns with the fact that the Confederate army, starting from scratch, enabled quicker promotions and a higher likelihood of becoming a general for people with officer training. It also had a higher ratio of generals to enlisted mean. On the other hand, joining the Union is associated with a lower risk of dying in the Civil War, reflecting the higher fatality rate on the Confederacy side, as reported in Column (3).

### v2

@@ Abstract
Do social networks and peer influence shape major life decisions in highly polarized settings? We explore this question by examining how peers influenced the allegiances of West Point cadets during the American Civil War. Leveraging quasi-random variations in the proportion of cadets from Free States, we analyze how these differences affected cadets’ decisions about which army to join. We have four main findings. First, there was a strong and significant peer effect: a higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. Second, the peer effect interacted with geography and economic circumstances: almost all cadets from Free States joined the Union Army (if they decided to join the war), and most cadets from Slave States that had more than a third of their population in slavery joined the Confederacy. The cadets who were most highly influenced were from Slave States that had less than a third of their population in slavery. Third, we analyze how cadets’ decisions affected their military rank and career outcomes. Fourth, we show that having served together in the Mexican-American war increased the peer influence, providing additional evidence that peer interaction was influential.
@@ 1 Introduction
Do social networks and peer influence shape individuals’ choices when societies are deeply polarized? This question is central not only to understanding moments of historical rupture but also to how political allegiances are formed in contemporary democracies. The literature on polarization has highlighted partisan sorting, media biases, divergent economic situations, and cultural divisions as key drivers of political alignment (e.g., Mason, 2018; Prior, 2013; Autor et al., 2020; Hochschild, 2016). Less is known about the role of peers and social networks in shaping political loyalties in polarized settings. While research has shown that peers matter for outcomes such as education, consumption, health, and careers, [^1] We cannot include a comprehensive review of the literature here, but such overviews are available in Benhabib, Bisin and Jackson (2010), Brechwald and Prinstein (2011), Aral and Nicolaides (2017), McGloin and Thomas (2019), Giletta et al. (2021), and Boucher et al. (2024). there is little causal evidence as to whether peers influence high-stakes political and/or life decisions, especially when these decisions come with severe personal risks and moral conflicts.
The American Civil War provides a unique opportunity to examine this question. West Point cadets decisions as to which side to fight for was often the most consequential choice of their lives, pitting personal backgrounds, economic interests, and regional loyalties against professional ties and peer influences. Although West Pointers played pivotal roles in the war, systematic empirical analysis of the factors influencing their allegiances remains limited. [^2] There exists a rich historical literature on West Pointers in the Civil War (e.g., Patterson, 2002). A recent political science study, White (2024), highlights the importance of economic interests. See also additional references below. This setting offers two key advantages. First, the quasi-random composition of cadet cohorts generates exogenous variation in exposure to peers from different regions. This allows us to separate peer effects from confounding factors (e.g., individuals typically self-select into like-minded groups) that typically make it difficult to disentangle network influence from shared beliefs or communal forces (see Manski, 1993; Angrist, 2014). Quasi-randomness arose from institutional features: although admissions broadly tracked the geographic distribution of congressional districts, annual recruitment varied substantially due to decentralized nomination procedures and qualification exams (e.g., Park, 1840). Second, West Point was the nation’s premier military academy, where cadets from Northern (Free) and Southern (Slave) lived and trained together, and forged close bonds before the war. This environment magnified the potential for peer influence. Historians have long noted that “group cohesion” and camaraderie motivated enlistment and fighting, but the evidence has been largely anecdotal, drawn from letters and diaries (e.g., McPherson, 1997, 2003; Siebold, 2007). Whether Northern peers swayed Southern cadets toward the Union, or whether local political and economic pressures outweighed such influence, has been an open question.
@@
To answer this, we digitize detailed biographical data on more than 1,600 West Point cadets from 1820 to 1860 and exploit quasi-random variation in peer composition. We classify states as Free or Slave based on whether slavery was legal in 1860 (corresponding to a slave population share exceeding 1%), and confirm robustness with alternative thresholds. The share of peers from Free States fluctuated annually without clear trends, creating plausibly exogenous exposure to Northern peers (see Section 3.1 for institutional details). This variation is not systematically correlated with individual backgrounds or home-state characteristics, whether in the full sample or when restricting to cadets who later fought in the war, supporting its interpretation as quasi-random. In the main text, we focus on the wartime choices of cadets who served in the conflict. The Supplemental Appendix presents additional robustness checks using the full cadet sample.
We have four main findings.
@@
The first is that there is significant and large peer effect on cadet’s decisions. A higher proportion of classmates from Free States significantly increased the likelihood that cadets from Slave States joined the Union Army. The effect is sizable: a one-standard deviation increase in the proportion of peers from Free States (an increase of roughly 5 out of 40 peers) raised the probability of joining the Union by four percentage points, or 11% of the mean. In contrast, nearly all cadets from Free States joined the Union regardless of cohort composition. The difference suggests a conflict between nationalism and sectionalism faced by cadets from Slave States, who were more susceptible to peer influence.
The second main finding is that the peer effect depends on background (economic) situations. Within the Slave States, we find that peer influence mattered little for cadets from states in which the slave population was high (e.g., above 1/3), but mattered strongly for cadets from Slave States with slave populations below 1/3, suggesting an interaction between peer influence and economic considerations rooted in the slave economy. Indeed, the peer influence was large and almost entirely concentrated among cadets from Border and Mid-level Slave States where support for secession was more divided.
@@
Our third main finding is that the cadets’ decisions had long-term implications for their careers and lives. Because many West Point cadets in our study later emerged as significant military figures, we can examine how their wartime choices influenced their post-war trajectories. Employing both OLS and an instrumental variable strategy (where home state and peer exposure predict allegiance), we find that for those from Slave States, those who sided with the Union experienced lower military ranks in 1865 and a diminished chance of becoming generals, but also a decreased likelihood of wartime mortality.
Our fourth finding examines the role of direct interactions in driving peer influence. We present two pieces of evidence. We find that peer influence is more pronounced within the same cohort than across cohorts, indicating the importance of direct, contemporaneous interactions. We also leverage variation in cadets’ likelihood of having co-fought in the Mexican-American War (1846–1848). We find that peer effects were stronger when slave-state cadets had more free-state peers with whom they had shared combat experience, and weaker when their co-combatants were also from Slave States.
We present several additional analyses to examine heterogeneity and test the robustness of our results, including tests that account for shifting political environments, peer dropouts, those not participating in the war, appointment states, and other factors.
Our study contributes knowledge about peer effects (see Benhabib, Bisin and Jackson (2010); Brechwald and Prinstein (2011); McGloin and Thomas (2019); Giletta et al. (2021) for some overviews), by demonstrating how peer influence can shape major life decisions, particularly during moments of extreme polarization. More broadly, it adds to the long-standing literature emphasizing the importance of “personal influence” in economic, political, and social behaviors (e.g., Katz and Lazarsfeld, 1955). Other research specifically involving soldiers concerns the spread of political ideas Jha and Wilkinson (2023); Grosjean et al. (2025), which our study complements in studying allegiance (and how it depends on economic circumstances).
@@
The experiences of West Point cadets during the American Civil War mirror other historical contexts where military academies produced leaders for both sides of major conflicts in divided nations. For instance, the Whampoa Military Academy in the 1920s produced commanders who fought on opposing sides during the Chinese Civil War (e.g., Bai, Jia and Wang, 2023), and the Heroic Military Academy in Mexico trained leaders who both supported and opposed the regime during the Mexican Revolution. Such decisions are not unique to military cadets, as similar choices are faced by ordinary citizens in many divided societies. However, analyzing peer influence in these broader contexts is often more challenging. The distinctive setting of the American Civil War, where allegiances were closely tied to geography, provides a unique opportunity to examine peer influence with a level of identification that is difficult to achieve in other cases.
@@
Finally, our study complements an expanding literature on leadership (e.g., Jones and Olken, 2005; Acemoglu and Jackson, 2015; Dippel and Heblich, 2021; Cagé et al., 2023; Bai, Jia and Yang, 2023; Ferrara, Dippel and Heblich, 2025). Understanding how leaders develop their ideas and loyalties is central to grasping important political and economic outcomes. Our findings highlight the significance of peer influence in shaping future leaders’ decisions during crises, demonstrating how peer influence can drive individual choices that may shape the course of history.
@@ 2.2 Geography and the Classification of States by High/Low Slave
Between 1820 and 1860, 1,656 cadets graduated from the United States Military Academy (see the complete records in Biographical Register of the Officers and Graduates of the United States Military Academy (Cullum, 1891)). 17 graduates were from outside the continental United States, and since there was only one graduate in 1813, with no peers, our analytical sample consists of 1,638 cadets. In our data, 964 and 674 cadets came from Free States and Slave States, respectively. Among those from Free States, 540 joined the armies and 91.9% of them joined the Union. Among those from Slave States, 389 joined the armies and 36.2% of them joined the Union.
Supplemental Appendix Table S2 presents the summary statistics of the key variables using the baseline threshold. Further, Supplemental Appendix Table S3 reports the cohort-level distribution of graduates by state type, along with their standard deviations, showing that there was substantial variation from cohort to cohort in the fraction of cadets from Free vs Slave States.
@@ 3.2 Statistical Analysis
$Union_{i,s,t}=\beta_{0}+\beta_{1}Peer_{i,t}+\gamma X_{i,s}+\alpha_{s}+\theta t+\epsilon_{i,s,t},$ (1)
where $Union_{i,s,t}$ denotes whether cadet $i$ from state $s$ and cohort $t$ joined the Union.
$Peer_{i,t}$ represents the proportion of peers from Free States. [^12] Note that our specification avoids the exclusion bias noted by Caeyers and Fafchamps (2024), since we are separating regressions for Slave and Free States and working with exogenous characteristics. $\alpha_{s}$ represents state fixed effects, whereas $X_{i,s}$ includes a range of personal and state characteristics such as the cadet’s age in 1860, their academic rank at West Point and the proportion of the slave population in their home state. Naturally, state characteristics are absorbed once home state fixed effects are accounted for.
We incorporate a linear time trend $t$ to account for potential temporal trends in Union enlistment. In addition, we report estimates that incorporate five-year bin fixed effects, which help absorb shifts in the political environment. To address potential serial correlation, we compute bootstrapped standard errors based on 400 resampling iterations. For robustness, we also re-estimate the model using standard errors clustered at the cohort level as well as bootstrapped standard errors clustered at the cohort level.
@@ 3.3 The Impact of Peers on Cadets’ Allegiance
Table 1 displays our main baseline results, indicating that peer influence matters for individuals from Slave States. Column (1) illustrates the raw correlation between the proportion of Free-State peers and the likelihood of joining the Union. Column (2) incorporates personal and state characteristics, while Column (3) further includes state fixed effects. To facilitate the interpretation of coefficients, we present the coefficient for a one standard deviation change in peer exposure. As shown, these estimates, between 0.040 and 0.062, are consistent across different models.
The standard errors are bootstrapped with 400 resampling iterations. In Supplemental Appendix Table S5, we also report standard errors clustered at the cohort level as well as those bootstrapped at the cohort level. The estimated coefficients on peer influence remain unchanged, while the corresponding p-values are generally smaller. To remain conservative, we rely on the larger standard errors in our main analysis.
The peer effect is sizable: a one standard deviation increase in of free state peers is roughly 5.5 more out of 40 cadets [^13] See Appendix Supplement Table S3. in the cohort being from Free States and leads to an 11 to 16% increase in the frequency of joining the union (mean 0.36). This is a conservative estimate. When using alternative thresholds for Slave States (5% and 10%), we find that peer effect raises the probability of Slave-State cadets joining the union by 17-19% (see Supplemental Appendix Table S6).
@@ 3.4 Career Outcomes
We know the post-war outcomes of many who joined an army, allowing us to investigate how their decisions during this critical historical juncture influenced their life results, such as military rank and survival probability. We do not presume that this had any influence on their allegiances, but does trace some consequences of their decisions.
As indicated in Columns (1)-(2) of Table 3, joining the Union correlates with a lower military rank in 1865 and a reduced probability of attaining a rank of General-level positions. This aligns with the fact that the Confederate army, was both smaller and had more generals per enlisted man, and starting from scratch had no existing officers, enabling quicker promotions and a higher likelihood of becoming a general for people with officer training. On the other hand, joining the Union is associated with a lower risk of dying in the Civil War, reflecting the higher fatality rate on the Confederacy side, as reported in Column (3).
@@ B. Slave-State cadets (1%-33%)
3.5 Peer Influence and Interaction
Our baseline results indicate that peers shaped individual decisions at a pivotal moment. To shed light on the mechanisms behind this influence, we focus on two complementary dimensions of interaction. First, we test whether peer effects are stronger within a cadet’s own cohort compared to across adjacent classes. Second, we assess whether the strength of influence depends on shared military experience: in particular joint participation in the Mexican-American War.
Influence by Cohort
We analyze peers by grouping them into cohorts and display the coefficients for peers within the same cohort, as well as for cohorts from t-1 to t-2 and t+1 to t+2, in Supplemental Appendix Figure S3. The coefficients in the figure are obtained from a single regression that includes all controls from our baseline analysis. The result reveals that peer influence is most important within the same cohort. In contrast, although there is some degree of influence from earlier cohorts (older cadets), future cohorts do not show comparable levels of influence (younger cadets).
Relevance of Co-fighting in the Mexican-American War
3.6 Additional Analyses
We conduct five supplementary analyses to account for (1) whether the effect of peers on cadets could vary with shifting political environments, (2) potential differences between peers who graduated and those who dropped out, (3) whether non-participants in the war affect the results, (4) whether cash-crop employment shapes peer influence, and (5) whether defining peers by appointment state rather than home state matters. We describe the results from these analyses below.
Shifting Political Environments
Although the variation in the composition of cadets is fairly random across time, there is a possibility that it interacts with changes in the political environment or other factors that could impact cadets’ choices. We conduct two analyses to account for such possibilities.
First, we divide the sample into periods before and after 1850. Although the issue of slavery was highly contentious throughout the period we study, it was particularly intense in the 1850s and nearly led to the collapse of the United States. In an effort to avert civil war, Congress passed the Compromise of 1850, which included the Fugitive Slave Law that obligated law enforcement nationwide to assist in capturing alleged runaway slaves, and many in the north resisted its enforcement (McPherson, 2003). Against this backdrop, we use 1850 as a dividing point to assess whether peer influence experienced notable changes before and after that year. As illustrated in Supplemental Appendix Table S10, peer influence (though estimated imprecisely due to limited sample sizes) persisted over time, with similar effects on either side of this divide.
Second, we augment our baseline specification with fixed effects at five-year intervals. This approach accounts for unobserved shocks or gradual shifts in the political environment that may coincide with changes in cadets’ decisions. As reported in Supplemental Appendix Table S11, the estimated effects remain highly consistent with our baseline results.
Taken together, these analyses indicate that, despite potentially evolving political circumstances, peer influence remained a significant factor in shaping cadets’ choices.
Graduate vs. Dropout Peer Influence
Including Those Who Were Not in the Armies
Considering Cash-Crop Employment
Considering Appointment State